OpenStudy (anonymous):

(n/6)^k-((n-1)/6)^k*(1-(n/6)^k) can I simplify this? I want to turn it into a sequence with a sigma notation but there's supposed to be a "cleaner" way of expressing it.

OpenStudy (anonymous):

\[(\frac{n}{6})^k-(\frac{n-1}{6})^k \times (1-(\frac{n}{6})^k)\]?

OpenStudy (anonymous):

exactly.

OpenStudy (anonymous):

actually wait.

OpenStudy (anonymous):

I should have probabily chosed a different letter for one of the ks, they're not always the same.

OpenStudy (anonymous):

the two ks on the left side of the multiplication sign can be different than the one on the right side.

OpenStudy (anonymous):

I cant edit it somehow i dont know whats wrong. But its the same, just (n/6)^q-((n-1)/6)^q*(1-(n/6)^k)